[figure: 9.7 cm]

EXCITONIC FERROMAGNETISM IN THE HEXABORRIDES

T.M. Rice Theor. Physik, ETH-Hoenggerberg, 8093 Zürich, Switzerland

Recently Young et al. [1] reported the surprising discovery of ferromagnetism in Ca_{1−x}La_{x}B_{6} - a material without d- or f-electrons. The ferromagnetism has special features: it occurs only in the narrow range 0 < x < 0.01, it has a very small moment(< 0.1 μ_{B}/La), but it has a high T_{C} (∼ 600K). The small overlap of valence and conduction bands provides favorable conditions for the formation of a triplet exciton condensate. When doped it is energetically favorable to distribute the carriers asymetrically between quasiparticle spin directions [2] leading to a ferromagnetic state with a small moment oriented perpendicular to the direction of the triplet condensate [3]. This form of excitonic ferromagnetism provides a natural explanation for the unusual properties of ferromagnetism in the hexaborides.

1. D.P. Young et al., Nature 397, 412 (1999)

2. B.A. Volkov,Yu. Kopaev and A.I. Rusinov, Sov.Phys. JETP 41, 952 (1975)

3. M.E. Zhitomirsky, T.M. Rice and V.I. Anisimov, cond-mat/9904330

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : T.M. Rice

Address for correspondence : Theor. Physik, ETH-Hoenggerberg, 8093 Zürich, Switzerland

Email address : rice@itp.phys.ethz.ch